OpenStudy (anonymous):

joseph wants to estimate 1 3/7 * 5 4/5 . To estimate, he simplifies the expression 1 1/2 * 6 . Is Joseph’s estimate a close estimate, an underestimate, or an overestimate? A. underestimate B. overestimate C. close estimate

OpenStudy (anonymous):

@DanJS @GretaKnows @Godlovesme

OpenStudy (danjs):

1 and a half is a rounded up 5 and 4/5 is rounded up it will be over the real value

OpenStudy (danjs):

he used two larger numbers than the original 2, they will end up having a product larger than the real actual product

OpenStudy (danjs):

(1+3/7)*(5+4/5) = 58/7 about 8.3 (1 + 1/2)*6 = 9 bigger than the real value

OpenStudy (danjs):

not sure how close, is close, for the close estimate